OpenStudy (agentnao):

At a given temperature, you have a mixture of benzene (vapor pressure of pure benzene = 745 torr) and toluene (vapor pressure of pure toluene = 290. torr). The mole fraction of benzene in the solution is 0.590. Assuming ideal behavior, calculate the mole fraction of toluene in the vapor above the solution.

OpenStudy (anonymous):

Use raoult's law, p=k x mole fraction of toluene = 1-0.59=0.31 p=290*0.31